OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (-7, 0) and a directrix at x = 7. @jim_thompson5910 I'm not sure what to do, can you help?

jimthompson5910 (jim_thompson5910):

what's the distance from the focus to the directrix?

OpenStudy (anonymous):

um i'm not sure. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

plot it out

jimthompson5910 (jim_thompson5910):

plot the line x = 7 and the point (-7,0)

OpenStudy (anonymous):

14?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

cut that in half to get 14/2 = 7

jimthompson5910 (jim_thompson5910):

so the focal distance is 7 units, which means p = 7

jimthompson5910 (jim_thompson5910):

the vertex will lie on the midpoint of (7,0) and (-7,0) what is the vertex?

OpenStudy (anonymous):

0,0?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

so (h,k) = (0,0)

jimthompson5910 (jim_thompson5910):

h = 0 k =0

jimthompson5910 (jim_thompson5910):

since the directrix is a vertical line, and because the focus is to the left of the directrix, we know the parabola opens up to the left

jimthompson5910 (jim_thompson5910):

so it's of the form -4p(x-h) = (y-k)^2 in this case p = 7 h = 0 k = 0 making our equation to be -4p(x-h) = (y-k)^2 -4*7(x-0) = (y-0)^2 -28x = y^2 y^2 = -28x

jimthompson5910 (jim_thompson5910):

you can also write the equation as y^2 + 28x = 0

OpenStudy (anonymous):

the only one i have with the anser y^2 is y2 = -14x the other answers are x = -(1/28)y^2 -28y = x2 y = -(1/28)x^2 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

hmm there has to be a typo then

jimthompson5910 (jim_thompson5910):

I don't see any other equivalent answers

jimthompson5910 (jim_thompson5910):

oh wait

jimthompson5910 (jim_thompson5910):

y^2 = -28x -28x = y^2 x = (-1/28)y^2

OpenStudy (anonymous):

Find an equation in standard form for the ellipse with the vertical major axis of length 16, and minor axis of length 4. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

the one you found

jimthompson5910 (jim_thompson5910):

what did I find?

OpenStudy (anonymous):

x = (-1/28)y^2

jimthompson5910 (jim_thompson5910):

that was the last problem

jimthompson5910 (jim_thompson5910):

and that's not even an ellipse

OpenStudy (anonymous):

yeah can you help me with this new one?

jimthompson5910 (jim_thompson5910):

what did you get for this new one

OpenStudy (anonymous):

i don't konw what to do?

jimthompson5910 (jim_thompson5910):

check your notes or go back to the lesson

OpenStudy (anonymous):

ok